Circular No. 6709

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1996 J1 (EVANS-DRINKWATER) Z. Sekanina, Jet Propulsion Laboratory and European Southern Ob-

servatory, writes: “A refined orbital solution (cf. IAUC 6662), based on my standard model for split comets (1982, in Comets, 251) and derived from 20 accurate separations of the secondary nucleus A from the primary B between 1997 May 9 and June 16 (MPC 29700 and 29967), suggests that the comet broke up on 1996 Oct. 11.1 ± 1.6 UT (or 80 days before perihelion) with a velocity of 0.93 ± 0.03 m/s in a direction nearly normal to the orbit plane, and that A drifted away from B with a relative nongravitational deceleration of 39.3 ± 0.3 units of 10^{−5} the solar attraction. Predicted separation distances and position angles for companion A relative to primary B (for 0^{h} UT on each date): 1997 July 31, 130′′, 268^{o}; Aug. 20, 146′′, 269^{o}; Sept. 9, 162′′, 269^{o}; Sept. 29, 177′′, 267^{o}; Oct. 19, 187′′, 265^{o}; Nov. 8, 187′′, 262^{o}; Nov. 28, 180′′, 260^{o}; Dec. 18, 168′′, 260^{o}.”

Total visual magnitude estimate by K. Hornoch, Lelekovice, Czech Republic (0.35-m reflector): July 13.98 UT, 13.6.

MARKARIAN 421 AND MARKARIAN 501 C. L. Bhat, Bhabha Atomic Research Centre (BARC), Mumbai, writes:

“The Imaging Element of the TeV γ-ray telescope array TACTIC (Bhat et al. 1994, Proc. ‘Towards a Major Atmospheric Cerenkov Detector III’, Univ. Acad., Tokyo) recently set up by BARC group at Mt. Abu, India, has been used to observe Mkn 501 and Mkn 421 in the tracking mode for ∼ 50 and 29 hr, respectively. The data have been analyzed using the ‘supercuts’ techniques (Reynolds et al. 1993, Ap.J. 404, 206). A 13.3σ confidence level γ-ray signal has been detected from Mkn 501 during April–May, with a time-averaged flux of (7.6 ± 0.7) × 10^{−11} photons cm^{−2} s^{−1} above 0.7 TeV (about twice the Crab flux). A peak flux of nearly four times that of the Crab has been recorded recently by the TACTIC on Apr. 13 (in agreement with the results reported by the Whipple, CAT, and HEGRA groups) and again on May 16. Significant flux variability has been observed on a day-to-day basis with pronounced dips on Apr. 12, May 14, and May 30. In the case of Mkn 421, our observations reveal that this source was in a comparatively low state of activity during Apr. 7–13 and May 7–29. A 3σ upper limit of 1.7 × 10^{−11} photons cm^{−2} s^{−1} above 0.8 TeV is deduced for the corresponding time-averaged flux.”

1997 July 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green